OpenStudy (breathlessreaper):

Can someone please help?? :< Which of the following statements best describes the effect of replacing the graph of y = f(x) with the graph of y = f(x) + 20? A. The graph of y = f(x) will shift left 20 units. B. The graph of y = f(x) will shift down 20 units. C. The graph of y = f(x) will shift up 20 units. D. The graph of y = f(x) will shift right 20 units

OpenStudy (anonymous):

I believe it is C o; because + 20 is on the y intercept that leaves you going either up or down because it is positive, it would be shifting up 20 units c:

OpenStudy (breathlessreaper):

Hmm o . o Okay Makes sense. Thank you ^^

OpenStudy (breathlessreaper):

Could you help with one more?

OpenStudy (anonymous):

I can try o:

OpenStudy (breathlessreaper):

For the function y = f(x), what is the ordered pair for the point on the graph when x = b - 2? A. (b - 2, f(b -2)) B. (x, f(b)) C. (x, b - 2) D. (b - 2, f(b))

OpenStudy (anonymous):

hmm i actually dont know o: im sorry :< but i wih you luck though ^~^

OpenStudy (anonymous):

wish*

OpenStudy (breathlessreaper):

Ill try to figure it out xD Thank you though.

OpenStudy (anonymous):

lol your welcome and thanks for the medal c:

jimthompson5910 (jim_thompson5910):

hint: if x = 2, then the point (2,f(2)) is on the graph of f(x) keep in mind that y = f(x)

OpenStudy (breathlessreaper):

@jim_thompson5910 Thank youuu

jimthompson5910 (jim_thompson5910):

you're welcome